The Surf's Up issues 1,000 shares of 6%, $100 par value preferred stock at the beginning of 2020. All remaining shares are common stock. The company was not able to pay dividends in 2020, but plans to pay dividends of $18,000 in 2021.

Assuming the preferred stock is cumulative, how much of the $18,000 dividend will be paid to preferred stockholders and how much will be paid to common stockholders in 2021?

Multiple Choice

  • $18,000 to preferred stockholders and $0 to common stockholders.

  • $6,000 to preferred stockholders and $12,000 to common stockholders.

  • $9,000 to preferred stockholders and $9,000 to common stockholders.

  • $12,000 to preferred stockholders and $6,000 to common stockholders.
Answer #1

Preferred dividends for 2020 = (1,000*100*6%) = 6,000

Preferred dividends for 2021 = (1,000*100*6%) = 6,000

Dividends to common stockholders = 18,000 - (6,000+6,000) = 6,000

Option D is the answer